# Artificial Intelligence LLM Large Language Model

### v2

@@ 3.5. Participants
A total of $N=526$ participants took part in the survey. A subsample of $n=116$ participants who regularly used both systems was included in the within-subject analyses. We determined our target sample size through an a priori power analysis ($f^{2}=0.05$, $\alpha=.05$, $\beta=.80$), which suggested approximately 250 participants for adequately powered within-subject analyses. The achieved subsample provides sufficient power to detect medium-sized effects, although smaller effects may remain undetected. Within the total sample, 186 participants completed the questionnaire for the university’s chatbot (customized LLMaaS) and 333 for ChatGPT (commercial LLM chatbot). An additional 123 participants reported using neither system at least once per month. Figure A1 shows an overview of participant groups according to their system usage.
@@ 3.6. Open Science
This study was preregistered on OSF [^3] Non-anonymous preregistration link removed for peer-review. Minor linguistic and structural adjustments were made to the preregistered hypotheses to improve clarity and consistency with the final study design. The dataset and supplementary materials (codebook, questionnaire overview) are available at OSF [^4] https://osf.io/vqsem/files/osfstorage?view_only=53ba0188a6db4a54aa08a1f34bf9078f. We excluded privacy-sensitive entries from the dataset prior to publication.
@@ 4.1. Descriptive Results: Chatbot Usage Patterns and Knowledge Test
Among the subsample of participants who had used both systems ($n=116$), the university chatbot was most commonly used several times a month ($32.8\%$) or several times a week ($25.0\%$), with only $9.5\%$ using it several times a day. In contrast, ChatGPT had a higher proportion of frequent users, with $19.0\%$ using it several times a day and $28.4\%$ several times a week (see Figure C1 for full distribution). Regarding usage purposes, both systems were most often used for reviewing or improving texts ($70.7\%$). Other prominent use cases included text creation ($54.3\%$ university chatbot; $61.2\%$ ChatGPT) and information search ($52.6\%$ university chatbot; $67.2\%$ ChatGPT; see Figure C2 for a more detailed overview). Awareness of model-switching functionality in the university chatbot was relatively high: $78.4\%$ of dual users knew about the option to change the underlying LLM. However, only $41.4\%$ had actively switched from the default model (GPT-4o) to another. Regarding performance on the knowledge test about the university’s customized LLMaaS, $24.1\%$ answered four questions correctly and $21.6\%$ achieved a perfect score. $14.7\%$ did not answer any item correctly (full distribution see Table C1).
@@ 4.2.1. Differences in Perceived Trust, Hallucinations, Privacy and Sustainability-Aware Use
We conducted paired t-tests (Table 3) to test H1 and H3-H6. To account for the increased risk of Type I error due to multiple hypotheses testing, we applied Holm’s step-down correction to the family of six statistical tests (Bender and Stefan Lange, 2001; Holm, 1979). This procedure controls the family-wise error rate (FWER) at $\alpha=0.05$ and was chosen because it is uniformly more powerful than the classical Bonferroni correction while still strictly controlling the FWER (Bender and Stefan Lange, 2001). The six $p$-values were ordered from smallest to largest and sequentially compared against adjusted α-thresholds (i. e., $\alpha$/$k$, where $k$ is the number of remaining tests). The procedure was terminated once a $p$-value exceeded the corresponding adjusted α-threshold; the respective hypothesis and all subsequent ones were rejected. H6, concerning sustainability-aware AI use, was the first hypothesis to exceed its adjusted α-threshold ($p=0.03$ vs. $\alpha=0.0167$), resulting in its rejection along with all subsequent hypotheses in the sequence. Consequently, H3, concerning cautious behavior toward hallucinations ($p=0.19$) and H2 (moderation: system type $\times$ trust in the university; $p=0.44$) were also rejected.
H1, H4 and H5 were supported, as the p-values remained below the respective adjusted $\alpha$-threshold: participants (H1) reported significantly higher levels of trust in the customized LLMaaS compared to ChatGPT, (H4) experienced significantly fewer hallucinations with the customized LLMaaS compared to ChatGPT, and (H5) reported significantly fewer privacy concerns for the customized LLMaaS compared to ChatGPT.
@@
Cautious behavior toward hallucinations (H3) 4.18 0.68 4.13 0.67 1.31 115 0.19 n/a4 0.12 [-0.06, 0.3]
@@
  • 1
    Holm-adjusted significance threshold per hypothesis rank (sequential correction).
  • 2
    Cohen’s d is reported as the standardized effect size for each comparison.
  • 3
    Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons.
  • 4
    Holm-adjusted thresholds are only assigned while p-values remain below α; correction stops after the first non-significant test.
  • 5
    Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.
4.2.2. Moderating Role of Organizational Trust
@@
Table 4. Mixed-Effects Model of system and trust in the university on user trust (number of observations: 232; $n=116$).
@@
TrustInUniversity_centered $272\textsuperscript{**} 0.0$840.084 $184.7$ $3.25$ $0.001$
System:TrustInUniversity_centered $0.066$ $0.085$ $114.0$ $0.78$ $0.440$
@@ 4.3. Controlling for Confounding Variables
4.3.1. Usage Frequency does not moderate the relationship between system type and experienced hallucinations
To test whether the difference in experienced hallucinations between the customized LLMaaS chatbot and ChatGPT was moderated by usage frequency, we estimated a Cumulative Logit Model (Table 5) including both main effects and their interaction using the VGAM package in R (Yee, 2010). Usage frequency did not significantly moderate the relationship between system type and experienced hallucinations, as indicated by a Wald test of all interaction terms ($\chi^{2}=9.60$, $df=5$, $p=0.09$). We chose this modeling approach because the usage frequency was measured with an ordinal scale. We accordingly transformed hallucination ratings into five ordered categories using equally spaced cut points. Among the usage frequency contrasts, the lowest p-value was observed for the fifth-order polynomial, which tests whether hallucination reports follow a complex pattern across usage levels. However, this effect did not fall below the alpha level of $0.05$ and was therefore not statistically significant.
Table 5. Cumulative Logit Model: System type and usage frequency on experienced hallucinations (number of observations: 227).
@@ 4.3.1. Usage Frequency does not moderate the relationship between system type and experienced hallucinations
$^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$
Model estimated on 227 observations (df = 893)
@@ 4.3.2. Model switching does not affect system differences
To examine whether switching the default language model (GPT-4o) within the university’s customized chatbot influenced the results, we repeated the within-subject comparisons using paired-samples t-tests on a reduced subsample of participants who reported never having changed the underlying model ($n=68$; see Appendix D for detailed test statistics). The pattern of results remained identical: the same differences between systems reached statistical significance (or not), and all observed effect directions were consistent with those in the full sample. Consequently, the interpretation of our findings and conclusions regarding the tested hypotheses remained unaffected. We therefore report the results based on the full sample of users who had experience with both systems ($n=116$) in the main text.
@@ 4.3.3. Prior knowledge does not affect system differences
To control for potential confounding effects of participants’ knowledge about the unversity’s customized LLMaaS chatbot, we included knowledge scores as a covariate in our analysis. The system differences proved robust to the prior knowledge of the participants, as neither the knowledge scores nor their interaction with the system type significantly predicted any outcome measures (all $p>0.11$; see Table D1 for detailed results). To examine this, we fitted Linear Mixed-Effects Models using the lme4 package, specifying fixed effects for system type, knowledge scores, and their interaction, as well as random intercepts for participants.
$4.$ Exploratory Analyses
@@
To complement the self-reported experienced hallucinations of users, we conducted two exploratory benchmark evaluations to assess the hallucination tendencies of the chatbots: TruthfulQA and HaluEva. The university’s customized LLMaaS chatbot and ChatGPT were tested using the same LLM, namely GPT-4o.
@@
TruthfulQA (Lin et al., 2022) is a benchmark designed to evaluate whether language models generate factually accurate responses to questions that are prone to false answers due to common misconceptions or false beliefs. A model performs well on this benchmark if it avoids reproducing widespread falsehoods learned from human-generated texts. We evaluated both systems on 50 questions from the TruthfulQA dataset. We selected 19 questions from the “Education & Science” category to reflect the university context of our study, and sampled 31 additional questions from the remaining categories. The university’s customized LLMaaS chatbot produced hallucinated answers in 25 out of 50 cases ($50\%$), whereas ChatGPT hallucinated in 19 out of 50 cases ($38\%$). Notably, the university chatbot hallucinated more frequently across both the domain-specific and general subsets of the benchmark.
@@ 4.4.3. Between-subjects comparisons mostly align with within-subjects results
To complement the within-subject analyses, we conducted exploratory between-subject comparisons using Welch’s two-sample t-tests. These analyses included participants who reported regular use of only one system – either the university’s customized LLMaaS chatbot ($n=70$) or ChatGPT ($n=217$). The full set of results is provided in Table E1; overall, the between-subjects results mostly aligned with the within-subjects results, and here we report only findings that deviated. Notably, participants who exclusively used the customized LLMaaS chatbot reported significantly higher sustainability-aware AI usage compared to those who only used ChatGPT. In contrast, the significant within-subject difference in trust in favor of the customized LLMaaS chatbot was not replicated in the between-subjects comparison. Similarly, the significant within-subject effect for perceived ease of use (favoring the customized LLMaaS chatbot) was not statistically significant in the between-subjects analysis. All other outcomes, including privacy concerns, experienced hallucinations, cautious behavior toward hallucinations and perceived usefulness, showed consistent patterns across both analysis approaches. Given that ChatGPT Pro users are not subject to token limits, which may compromise the comparability of groups for sustainability-aware AI usage, we conducted an additional comparison between exclusive ChatGPT Free users ($n=176$) and users of the customized LLMaaS chatbot ($n=70$). Detailed results are provided in Table E2. This analysis showed the same pattern of results as the full-sample between-subjects comparison, including significantly higher sustainability-aware AI usage among customized LLMaaS users and no significant difference in trust.
@@
Participants were instructed to respond to all dependent variable items with regard to their study- or work-related use of the AI systems in the context of their everyday university life. However,we also report on the overall usage tendencies that participants indicated across the private vs. work/study usage context. Participants who used both systems reported using the customized LLMaaS chatbot with a strong orientation toward study- or work-related contexts ($M=5.48$, $SD=0.81$; scale range: 161-6), whereas reported usage of ChatGPT was more balanced across the spectrum ($M=3.32$, $SD=1.61$). A paired samples t-test confirmed that this difference was statistically significant, $t(115)=13.36$, $p<.001$, $d=1.24$. Notably, $64.7\%$ of participants selected the highest possible work/study-related option (6) for the customized LLMaaS chatbot, while only $11.2\%$ did so for ChatGPT.
@@ 5. Discussion
This study examined how customizations of LLMaaS chatbots affect users’ trust, perception and usage of such tools in the context of everyday university life. Our findings support the hypothesized differences in users’ perceptions of the customized LLMaaS chatbot versus ChatGPT in regard to trust, privacy and experienced hallucinations. In the following, we interpret the results and elaborate on challenges in achieving calibrated user perceptions of LLMaaS. Finally, we discuss the theoretical and practical implications, as well as the limitations, of our study.
@@
Participants showed higher trust in the university’s customized LLMaaS chatbot compared to ChatGPT (H1). This finding supports the assumption that users draw on cues about the system such as the corporate interface design in their trust assessments. It also aligns with the underlying theory of trust in automation (Lee and See, 2004), which highlights process and purpose as key dimensions of user trust. Our findings support the relevance of these dimensions in shaping trust in LLM chatbots, as the customized LLMaaS chatbot may have been perceived as more easily accessible and integrated into institutional structures (process), and as provided by a benevolent, responsible actor (purpose), thereby fostering greater trust. The result further supports our rationale that customizations may enhance user trust by increasing factors such as perceived openness or transparency of the system, which is consistent with previous findings (Schwartz et al., 2023; Nordheim et al., 2019). Moreover, it suggests that the role of familiarity in fostering interpersonal trust (Alarcon et al., 2016) might also be relevant in human–chatbot interactions.
The hypothesized moderation effect of trust in the university (H2) was not supported, i. e., the effect of system type on user trust did not depend on how much users trusted the university. Instead, both the system type (customized LLMaaS chatbot) and higher trust in the university independently contributed to higher trust in the system. This finding may reflect the influence of users’ general propensity to trust, that is, a dispositional tendency to attribute trustworthiness across different social and technological entities (Mcknight et al., 2011; Heyns and Rothmann, 2015). Overall, the results point to additive rather than interactive effects of system type and trust in the university on users trust in the system.
We found no statistically significant difference in how cautious participants behaved toward hallucinations between the customized LLMaaS chatbot and ChatGPT (H3). However, participants reported more experienced hallucinations with ChatGPT than with the customized LLMaaS chatbot (H4). This combination of results is noteworthy as we expected that the same customizations of the LLMaaS chatbot, the absence of an in-chat hallucination warning and the corporate design, would jointly reduce users’ alertness toward hallucinations, thereby leading to both less cautious behavior and fewer reported hallucinations. However, our findings only support the latter. While the lack of an effect for H3 could be due to the low internal consistency of our self-developed scale for cautious behavior, we discuss further possible interpretations. One possible explanation for this pattern is that interface customizations may have reduced users’ mental alertness only to a limited degree. This may have influenced the detection of hallucinations, which rely more on automatic or intuitive processes. However, it may not have affected more reflective, effortful strategies for preventing hallucinations or verifying outputs. These interpretations gain further nuance when considered alongside our exploratory benchmark evaluations. TruthfulQA directly assessed hallucination generation and indicated that the LLMaaS chatbot in fact produced more hallucinated answers than ChatGPT. In contrast, HaluEval evaluated hallucination detection performance and showed that the university chatbot slightly outperformed ChatGPT. Taken together, these results suggest a miscalibration between users’ perceptions and objective performance: although participants perceived the customized chatbot as less hallucination-prone, benchmark evidence indicates that it may generate hallucinations more frequently, while at the same time showing somewhat stronger abilities in recognizing them. This discrepancy underscores the role of trust-inducing cues and institutional framing in shaping users’ vigilance toward hallucinations, potentially leading to an underestimation of risks.
@@
Higher privacy perceptions were observed for the customized LLMaaS chatbot compared to ChatGPT (H5). These results relate to actual system differences in terms of privacy safeguards (see Figure 1). Additionally, users might have associated the university-hosted system with more responsible or secure data practices due to its institutional branding in the interface design or login-based access within the university network. Prior work has shown that strong corporate branding can reduce perceived privacy risks and increase users’ willingness to disclose personal information online (Wang, 2019; Myerscough et al., 2008; Jain et al., 2022). In line with this reasoning, the university’s institutional branding may have served as a trust signal, e. g., leading users to perceive the system as more accountable and less likely to monetize their data. These findings highlight the role of contextual design cues in shaping perceived privacy and suggest that institutional trust signals may function as privacy heuristics in AI-mediated interaction.
@@
Participants did not report more sustainability-aware AI usage when interacting with the customized LLMaaS chatbot compared to ChatGPT (H6). This finding stands in contrast to previous work showing that making resource consumption visible can foster more sustainable behavior (Penkert and Gröschel, 2023; Sanduleac et al., 2017). Notably, the exploratory between-subjects comparisons showed different results: users who exclusively used the customized LLMaaS chatbot reported more sustainability-aware AI usage than those who exclusively used ChatGPT (this finding also held true when comparing LLMaaS-exclusive users to users of the free, token-restricted ChatGPT version). Interestingly, LLMaaS-exclusive users also reported higher sustainability-aware AI usage than the group of users of both systems showed for either the LLMaaS chatbot or ChatGPT. This pattern could reflect underlying differences between user groups, such that individuals who use only the customized LLMaaS chatbot may differ systematically in values (Shang et al., 2023) or usage motivations (Wunderlich et al., 2019) that relate to more sustainability-aware AI behavior. Alternatively, lower sustainability-aware AI usage associated with ChatGPT might originate from characteristics of the system itself (e. g., no token usage visualization; low-barrier access across devices, including as a mobile app) and carry over to other systems used by the same person. While these interpretations remain speculative, they highlight a promising avenue for future research on how individual traits or system characteristics may shape sustainability-aware AI usage.
Overall, our results support the assumption that customization-related process and purpose cues enhance user trust, and that institutional framing contributes to privacy perceptions. However, they also raise questions about whether such cues affect deliberate behaviors, such as cautious behavior toward hallucinations. The lack of moderation by institutional trust and the mixed results on sustainability-aware AI use suggest that, alongside customization effects, individual traits – such as the propensity to trust or sustainability-related values – might represent a complementary factor that is worth examining in future work on human–AI interaction.
5.2. How Calibrated Are Participants’ Perceptions of the Chatbots?
The concept of calibration has previously been considered primarily in relation to trust (Parasuraman and Riley, 1997; Lee and See, 2004). In this study, we conceptualize the concept of calibration more broadly and build on the TrAM model (Schlicker et al., 2025) by linking actual system features with perceived system characteristics via cues experienced by the user. In general, we see that users displayed higher trust, perceived greater privacy and experienced less hallucinations with the customized LLMaaS chatbot. At first glance, this suggests that users are able to calibrate their perceptions towards an LLM chatbot, as, e. g., the university’s customized chatbot has stronger privacy safeguards and users have less concerns. However, we question whether the extent of the perceived difference is justified and if users’ consequent behavior is “right”. First, we question how well users understand the differences between the systems, given the results of the knowledge test (see Section 4.3). Second, an appropriate calibrated perception is further questioned by the results of the hallucination benchmark evaluations, which allowed us to relate users’ subjective experiences to objective performance metrics under standardized conditions. The results showed that, although participants perceived fewer hallucinations with the customized LLMaaS, objective metrics suggest that this system was at least equally, if not more, prone to hallucinations. Accordingly, the anticipated link between the system’s lower temperature setting and reduced hallucination tendency (see Figure 1) did not hold. In sum, this divergence highlights a critical gap in users’ hallucination alertness calibration, that is, their ability to align vigilance against hallucinations with the system’s actual risk. Users may underestimate the risk of hallucinations due to trust-inducing cues such as branding, institutional context, and lack of persistent warnings, potentially compromising critical engagement with LLM outputs. When this finding is considered alongside the significantly higher perceived privacy in LLMaaS, the question arises as to whether users are overtrusting the customized LLMaaS chatbot.
The results of this study advance the concept of calibrated trust by extending it to broader calibrated system perceptions, including privacy perception calibration and hallucination alertness calibration. By integrating the TrAM framework, we have demonstrated the importance of providing users with accurate system information, particularly in relation to customizations, via cues to enable calibrated system perceptions that support safe and appropriate use.

### v3

@@ 3.5. Participants
Participants were recruited via the university chatbot landing page, departmental communications, campus flyers, the university news website, and official university social media channels. A total of $N=526$ participants took part in the survey. $n=116$ participants used both systems regularly. We determined our target sample size through a priori power analysis ($f^{2}=0.05$, $\alpha=.05$, $\beta=.80$), which suggested N = 250 for within-subject analyses. The achieved subsample provides sufficient power to detect medium-sized effects, although smaller effects may remain undetected. Within the total sample, 186 participants completed the questionnaire for the university’s chatbot (customized LLMaaS) and 333 for ChatGPT (commercial LLM chatbot). An additional 123 participants reported using neither system at least once per month. Figure A1 shows an overview of participant groups according to their system usage.
@@ 3.6. Open Science
This study was preregistered on OSF [^3] https://osf.io/kjcg5/overview?view_only=2e3c090e4d294fe8b714550a714bcb1e. Minor linguistic and structural adjustments were made to the preregistered hypotheses to improve clarity and consistency with the final study design. The dataset and supplementary materials (codebook, questionnaire overview) are available at OSF [^4] https://osf.io/vqsem/files/osfstorage?view_only=53ba0188a6db4a54aa08a1f34bf9078f. We excluded privacy-sensitive entries from the dataset prior to publication.
@@ 4.1. Descriptive Results: Chatbot Usage Patterns and Knowledge Test
Among the subsample of participants who had used both systems ($n=116$), the university chatbot was most commonly used several times a month ($32.8\%$) or several times a week ($25.0\%$). In contrast, ChatGPT had a higher proportion of frequent users, with $19.0\%$ using it several times a day and $28.4\%$ several times a week (see Figure C1 for full distribution). Regarding usage purposes, both systems were most often used for reviewing or improving texts ($70.7\%$), text creation ($54.3\%$ university chatbot; $61.2\%$ ChatGPT) and information search ($52.6\%$ university chatbot; $67.2\%$ ChatGPT; see Figure C2). Awareness of model-switching functionality in the university chatbot was relatively high: $78.4\%$ of dual users knew about this option. However, only $41.4\%$ had actively switched from the default model (GPT-4o). Regarding performance on the knowledge test about the university’s customized LLMaaS, $24.1\%$ answered four questions correctly and $21.6\%$ achieved a perfect score. $14.7\%$ did not answer any item correctly (see Table C1).
@@ 4.2.1. Differences in Perceived Trust, Hallucinations, Privacy and Sustainability-Aware Use
We conducted paired t-tests (Table 2) to test H1 and H3-H6. To account for the increased risk of Type I error due to multiple hypotheses testing, we applied Holm’s step-down correction to the family of six statistical tests (Bender and Stefan Lange, 2001; Holm, 1979). This procedure controls the family-wise error rate (FWER) at $\alpha=0.05$ and was chosen because it is uniformly more powerful than the classical Bonferroni correction while still strictly controlling the FWER (Bender and Stefan Lange, 2001). The six $p$-values were ordered from smallest to largest and sequentially compared against adjusted α-thresholds (i. e., $\alpha$/$k$, where $k$ is the number of remaining tests). The procedure was terminated once a $p$-value exceeded the corresponding adjusted α-threshold; the respective hypothesis and all subsequent ones were rejected. H6, concerning sustainability-aware AI use, was the first hypothesis to exceed its adjusted α-threshold ($p=0.03$ vs. $\alpha=0.0167$), resulting in its rejection along with all subsequent hypotheses. Consequently, H3, concerning cautious behavior toward hallucinations ($p=0.19$) and H2 (moderation: system type $\times$ trust in the university; $p=0.44$) were also rejected.
H1, H4 and H5 were supported, as the p-values remained below the respective adjusted $\alpha$-threshold: participants reported significantly (H1) higher levels of trust, (H4) fewer perceived hallucinations and (H5) fewer privacy concerns for the customized LLMaaS compared to ChatGPT.
@@
Cautious behavior toward hallucinations (H3) 4.18 0.68 4.13 0.67 1.31 115 0.19 n/a4 0.12 [-0.06, 0.30]
@@
Notes. 1 Holm-adjusted significance threshold per hypothesis rank (sequential correction). 2 Cohen’s d is reported as the standardized effect size for each comparison. 3 Confidence intervals (CI) for Cohen’s d are not adjusted for multiple comparisons. 4 Holm-adjusted thresholds are only assigned while p-values remain below $\alpha$; correction stops after the first non-significant test. 5 Ease of Use and Usefulness were analyzed exploratively and were not part of the Holm correction.
4.2.2. Moderating Role of Trust in the University
@@
Table 3. Mixed-Effects Model of system and trust in the university on user trust (number of observations: 232; $\textbf{n}=\textbf{116}$).
@@
TrustUniversity_centered $272\textsuperscript{**} 0.0$840.084 $184.7$ $3.25$ $0.001$
System:TrustUniversity_centered $0.066$ $0.085$ $114.0$ $0.78$ $0.440$
@@ 4.3. Controlling for Confounding Variables
4.3.1. Usage Frequency does not moderate the relationship between system type and perceived hallucinations
To test whether the difference in perceived hallucinations between the customized LLMaaS chatbot and ChatGPT was moderated by usage frequency, we estimated a Cumulative Logit Model (Table 4) including both main effects and their interaction using the VGAM package in R (Yee, 2010). Usage frequency did not significantly moderate the relationship between system type and perceived hallucinations, as indicated by a Wald test of all interaction terms ($\chi^{2}=9.60$, $df=5$, $p=0.09$). We chose this modeling approach because the usage frequency was measured with an ordinal scale. We accordingly transformed hallucination ratings into five ordered categories using equally spaced cut-off points. Among the usage frequency contrasts, the lowest p-value was observed for the fifth-order polynomial, which tests whether hallucination reports follow a complex pattern across usage levels. However, this effect did not fall below the alpha level of $0.05$ and was therefore not statistically significant.
Table 4. Cumulative Logit Model: System type and usage frequency on perceived hallucinations (number of observations: 227).
@@ 4.3.1. Usage Frequency does not moderate the relationship between system type and perceived hallucinations
Notes. $^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$.
Model estimated on 227 observations (df = 893).
@@ 4.3.2. Model switching does not affect system differences
To examine whether switching the default language model (GPT-4o) within the university’s customized chatbot influenced the results, we repeated the within-subject comparisons using paired-samples t-tests on a reduced subsample of participants who reported never having changed the underlying model ($n=68$; see Table D1 for detailed test statistics). The pattern of results remained identical: the same differences between systems reached statistical significance (or not), and all observed effect directions were consistent with those in the full sample.
@@ 4.3.3. Prior knowledge does not affect system differences
To control for potential confounding effects of participants’ knowledge about the university’s customized LLMaaS chatbot, we included knowledge scores as a covariate (using Linear Mixed-Effects Models with the lme4 package, with fixed effects for system type, knowledge scores, and their interaction). The system differences proved robust to the prior knowledge of the participants (all $p>0.11$; see Table D2).
4.4. Assessment of Common-Method Bias
Because all key constructs were measured via self-report, we examined the risk of common-method bias. We focused on analyses that modeled associations between multiple predictors (i. e., the mixed-effects moderation model in Section 4.2.2 and the ordinal regression in Section 4.3.1), as within-subject mean comparisons are largely unaffected because any method factor cancels out in the difference scores. Following the full-collinearity approach, we computed variance inflation factors (VIFs) from equivalent linear models including all fixed effects (Kock, 2015). In the mixed-effects model, all predictors (system type, trust in the university, and their interaction) showed VIFs $\leq 2.0$ (see Table E1). In the ordinal model, adjusted generalized VIFs $\bigl(\operatorname{GVIF}^{1/(2\cdot df)}\bigr)$ (Fox and Monette, 1992) for system type, usage frequency, and their interaction were $\leq 2.67$ (see Table E2). These values are well below conservative thresholds (3.3; (Cenfetelli and Bassellier, 2009; Petter et al., 2007)) and indicate no problematic multicollinearity, nor evidence of a dominant common-method factor.
4.5. Exploratory Analyses
@@
To complement the self-reported perceived hallucinations of users, we conducted two exploratory benchmark evaluations to assess the hallucination tendencies of the chatbots: TruthfulQA and HaluEval. The university’s customized LLMaaS chatbot and ChatGPT were tested using the same LLM, namely GPT-4o.
@@
TruthfulQA (Lin et al., 2022) is a benchmark designed to evaluate whether language models generate factually accurate responses to questions that are prone to false answers due to common misconceptions. A model performs well if it avoids reproducing widespread falsehoods. We evaluated both systems on 50 questions from the TruthfulQA dataset including 19 questions from the “Education & Science” category to reflect the university context of our study and 31 questions from the remaining categories. The university’s customized LLMaaS chatbot produced hallucinated answers in 25 out of 50 cases ($50\%$), whereas ChatGPT hallucinated in 19 out of 50 cases ($38\%$). Notably, the university chatbot hallucinated more frequently across both the domain-specific and general subsets of the benchmark.
@@ 4.5.3. Between-subjects comparisons mostly align with within-subjects results
To complement the within-subject analyses, we conducted exploratory between-subject comparisons using Welch’s two-sample t-tests. These analyses included participants who reported regular use of only one system – either the university’s customized LLMaaS chatbot ($n=70$) or ChatGPT ($n=217$). The full set of results is provided in Table F1; overall, the between-subjects results mostly aligned with the within-subjects results, and here we report only findings that deviated. Participants who exclusively used the customized LLMaaS chatbot reported significantly higher sustainability-aware AI usage compared to those who only used ChatGPT. The significant within-subject difference in trust in favor of the customized LLMaaS chatbot was not replicated. Similarly, the significant within-subject effect for perceived ease of use (favoring the customized LLMaaS chatbot) was not statistically significant. Given that ChatGPT Pro users are not subject to token limits, which may compromise the comparability of groups for sustainability-aware AI usage, we conducted an additional comparison between exclusive ChatGPT Free users ($n=176$) and users of the customized LLMaaS chatbot ($n=70$). The same pattern of results were obtained as with the full-sample (see Table F2).
@@
Participants were instructed to answer items with regard to their study- or work-related use of the AI systems. However, we also asked about the overall usage tendencies. Participants who used both systems reported using the customized LLMaaS chatbot with a strong orientation toward study- or work-related contexts ($M=5.48$, $SD=0.81$; scale range: 161-6), whereas reported usage of ChatGPT was more balanced across the spectrum ($M=3.32$, $SD=1.61$). A paired samples t-test confirmed that this difference was statistically significant, $t(115)=13.36$, $p<.001$, $d=1.24$.
@@ 5. Discussion
This study examined how customized LLMaaS chatbots are perceived and used differently compared to commercially available alternatives in the context of everyday university life. Our findings support the hypothesized differences in users’ perceptions in regard to trust, privacy and perceived hallucinations.
@@
The findings demonstrate that students and staff perceive the customized LLMaaS chatbot differently compared to ChatGPT. This is a significant finding, as educational institutions are increasingly making high investments in generative AI solutions (Uspenskyi, 2025), and there has been no comparative study of the perception of such solutions from the user’s perspective to date.
Participants showed higher trust (H1) and fewer privacy concerns (H5) with regards to the university’s customized LLMaaS chatbot compared to ChatGPT
@@
It is conceivable that the differences between the systems enhance user trust and perceived privacy by increasing factors such as perceived openness or transparency of the system, which is consistent with previous findings (Schwartz et al., 2023; Nordheim et al., 2019). Additionally, users might have associated the university-hosted system with more responsible or secure data practices due to its institutional branding in the interface design or login-based access within the university network. This adds to general work from marketing research that showed that strong corporate branding can reduce perceived privacy risks and increase users’ willingness to disclose personal information online (Wang, 2019; Myerscough et al., 2008; Jain et al., 2022).
The hypothesized moderation effect of trust in the university (H2) was not supported.
The effect of system type on user trust did not depend on how much users trusted the university. Instead, both the system type (customized LLMaaS chatbot) and higher trust in the university independently contributed to higher trust in the system. This finding may reflect the influence of users’ general propensity to trust, that is, a dispositional tendency to attribute trustworthiness across different social and technological entities (Mcknight et al., 2011; Heyns and Rothmann, 2015). Overall, the results point to additive rather than interactive effects of system type and trust in the university on users’ trust in the system.
Participants did not behave more cautiously toward hallucinations (H3), but reported more perceived hallucinations with ChatGPT than with the customized LLMaaS chatbot (H4).
This combination of results is noteworthy as we expected that the LLMaaS chatbot would jointly be associated with less cautious behavior and fewer perceived hallucinations compared to ChatGPT, both resulting from lower user alertness toward hallucinations. However, our findings only support fewer perceived hallucinations (H4). While the lack of an effect for H3 could be due to the low internal consistency of our self-developed scale, there is another possible explanation: The interface customizations may have reduced users’ mental alertness only to a limited degree, influencing the detection of hallucinations, which rely more on automatic or intuitive processes, but not more reflective, effortful strategies for preventing hallucinations or verifying outputs.
Concentrating on H4, this finding is particularly interesting when considered alongside our exploratory benchmark evaluations. TruthfulQA directly assessed hallucination generation and indicated that the LLMaaS chatbot actually produced more hallucinated answers than ChatGPT. Because the customized LLMaaS chatbot operated at a lower temperature than ChatGPT, one might expect it to hallucinate less. Yet, this benchmark suggests that the temperature difference did not reduce hallucination generation and therefore cannot account for the lower level of perceived hallucinations reported by participants. HaluEval evaluated hallucination detection performance and showed a minor difference, with the university chatbot outperforming ChatGPT. Taken together, these results suggest a miscalibration between users’ perceptions and objective performance: participants apparently notice less hallucinations in the customized LLMaaS chatbot, although benchmark evidence indicates that it may generate hallucinations more frequently. This discrepancy may indicate that the institutional framing and the absence of an in-chat warning could have reduced users’ attentiveness to hallucinations, thereby contributing to an underestimation of risks.
Participants did not report more sustainability-aware AI usage when interacting with the customized LLMaaS chatbot compared to ChatGPT (H6).
@@
This finding stands in contrast to previous work showing that making resource consumption visible can foster more sustainable behavior (Penkert and Gröschel, 2023; Sanduleac et al., 2017). The exploratory between-subjects comparisons showed that users who exclusively used the customized LLMaaS chatbot reported more sustainability-aware AI usage than those who exclusively used ChatGPT. This could reflect underlying differences between user groups, e.g., regarding their values (Shang et al., 2023) or usage motivations (Wunderlich et al., 2019). Alternatively, lower sustainability-aware AI usage associated with ChatGPT might originate from characteristics of the system itself (e. g., no token usage visualization; low-barrier access across devices). The results highlight a promising avenue for future research on how individual traits or system characteristics may shape sustainability-aware AI usage.